OpenStudy (anonymous):

A sign designer created a sign in the shape shown below. Shannon makes large cookies. These cookies are 22 inches in diameter. How much icing will she need to cover the top of each cookie?

OpenStudy (ghazi):

area of icing that is needed to coverup is = 66*40

OpenStudy (anonymous):

121π in2 484π in2 11π in2 22π in2

OpenStudy (ghazi):

i guess it's 484 pi inch square

OpenStudy (ghazi):

total area= 66*40 - (3*pi * (11)^2)= did you get this?

OpenStudy (anonymous):

i confused myself n then realized it was two questions in one. I am very sorry for the confusion but imma give u a medal anyways. sorry for confusion

OpenStudy (ghazi):

it's easy.....see there are 6 semi circle making it 3 circles of radius 11 (since diameter is 22) and total area of it = 66*40 now you've to subtract area of 6 semi circle which gives total area of icing = 66*40-(3*3.14*11*11) =484 pi